OpenStudy (anonymous):

to find the intercepts of: 6x-2y=12, i have to simplify the entire equation right?

zepdrix (zepdrix):

Yes! :) Let's get the equation into slope-intercept form first. y=mx+b So we want to get y by it's lonely self. Do you need some help with that step? :)

OpenStudy (anonymous):

Constant-divided-by-coefficient works well too.

OpenStudy (anonymous):

Slope-intercept form is not necessary to find intercepts.

OpenStudy (anonymous):

so it would be -2y=-6+12 y=-3-6 ?

OpenStudy (anonymous):

(It's still good to have anyway . . .) Oops, you lost an x there.

zepdrix (zepdrix):

Woops! Your x fell out somewhere :O

zepdrix (zepdrix):

What's wrong with slope-intercept form cliff? :| it's much to use than that ugly point slope form...

zepdrix (zepdrix):

much nicer*

OpenStudy (anonymous):

y=-3x-6 ?

OpenStudy (anonymous):

OP is looking for intercepts, not a graph.

zepdrix (zepdrix):

Yay gj briana c:

OpenStudy (anonymous):

so how do i find the intercepts?

zepdrix (zepdrix):

woops, i think the x is positive c: Make a little boo boo in there? Hmm

OpenStudy (anonymous):

In general, though, if you have a linear equation in standard form, Ax+By=C, the x-intercept is given by C/A, and the y-intercept is C/B. This is equivalent to setting y=0 and solving for x to get x-intercept, and then setting x=0 and solving for y to get y-intercept, except it gets both intercepts simultaneously with no manipulation of the equation whatsoever.

zepdrix (zepdrix):

Oh i see c:

OpenStudy (anonymous):

I figured, since that's what Brianna asked for, might as well go directly to it. Slope-intercept form is still great for graphing, and I recommend doing that too to verify the intercepts pictorially.

OpenStudy (anonymous):

so its (2,0) and (0,6) ?

OpenStudy (anonymous):

(as an aside, another handy thing to do sometimes is to divide every term in the equation by their greatest common factor to get an equivalent equation with smaller coefficients. i.e. 6x-2y=12 ÷ 2 -> 3x-y=6. This new equation is equivalent because it has the same solutions.

OpenStudy (anonymous):

Don't forget to include the negative sign with the coefficient.

OpenStudy (anonymous):

The y-intercept would be 12 ÷ -2 (or 6 ÷ -1) = -6

OpenStudy (anonymous):

(0,-6) ?

OpenStudy (anonymous):

Now you got it. Try graphing that line to see that it actually does go through those intercepts.